# latexpage numbers missing

### v1

@@ Abstract
Keywords: amortized inference, approximate simulation, diffusion models, generative models, likelihood-free inference, simulation-based inference, spatial extremes, uncertainty quantification
@@ 1 Introduction
In recent years, neural networks have been increasingly used to tackle computational challenges in the first step of data analysis: model fitting through parameter inference (Louppe et al., 2016; Cranmer et al., 2020; Dalmasso et al., 2020, 2024). These techniques have been applied in various ways to the spatial setting. For instance, Walchessen et al. (2024) leverage neural networks to efficiently evaluate the likelihood function of a spatial process model, while Lenzi et al. (2023) and Sainsbury-Dale et al. (2024, 2025) employ neural networks as Bayes estimators for parameter estimation. Both approaches rely on simulation-based inference (SBI) paired with machine learning—the practice of utilizing simulations from a statistical model to train a machine learning model for statistical inference. These methods preclude the need for an explicitly known or computationally tractable likelihood function (Cranmer et al., 2020; Dalmasso et al., 2020). In contrast, less work has focused on the subsequent—yet equally important—spatial inference task: making predictions and quantifying uncertainty at unobserved locations when the predictive distributions are not available in closed form. Existing methods, which typically do not use SBI and machine learning, include Markov chain Monte Carlo (MCMC) (Gelfand et al., 2010) and variational inference (Jordan et al., 1999; Banerjee et al., 2014). However, these approaches often involve a trade-off between computational efficiency and statistical accuracy and usually become impractical when dealing with a large number of observations or prediction locations.
@@
Here, we propose a new method using SBI to train a neural generative model that can simulate from predictive distributions in a spatial setting. Specifically, we employ state-of-the-art diffusion models that are used for image generation (Yang et al., 2023), because regularly gridded spatial data can effectively be viewed as image data. These models leverage both forward diffusion and backward evolution to generate samples from a target distribution in a computationally efficient manner. In the forward diffusion process, samples from a complex target distribution are progressively transformed into samples from a simpler reference distribution. This forward process, described by a stochastic differential equation (SDE), is designed to ensure the existence of a corresponding reverse SDE that evolves samples from the reference distribution back into samples from the target distribution (Song et al., 2021). This reverse SDE is characterized by a combination of user-defined terms from the forward SDE and an unknown score function, which is approximated using a neural network trained via a specifically designed loss function. Once the neural network is trained, a sample from the complex target distribution can be drawn efficiently by simulating from the simple reference distribution and applying the learned reversed SDE.
@@
There are several approaches to incorporating conditional information into a diffusion process, many of which originate from the inpainting problem, a topic extensively studied in computer vision (e.g., Lugmayr et al., 2022; Quan et al., 2024; Simkus and Gutmann, 2025). Generally, in image inpainting, the objective is to produce realistic, yet diverse images from a partial image. These approaches usually make no assumptions about an underlying statistical model and, consequently, are not designed to approximate any true predictive distributions. In contrast, Cardoso and Pereira (2025) adopt an approach that first trains a diffusion model to unconditionally sample from a spatial process and then performs conditional simulation from the predictive distribution via training-free guidance. However, this approach has not been applied to non-Gaussian processes and relies on several approximations during conditional simulation, which can lead to slower and potentially less accurate generation. A few inpainting methods, such as twisted diffusion (Wu et al., 2023) and controllable generation (Song et al., 2021), have asymptotic guarantees for exact conditional simulation. Yet, these methods have not thus far been validated in a statistically rigorous manner and may not scale easily to high-dimensional data.
@@
A recent contribution by Gloeckler et al. (2024) provides a masking approach to incorporate conditional information into the score-based diffusion process that is designed to accurately simulate from conditional distributions of parameters and data including posterior distributions. However, their approach was not designed for high-dimensional conditioning, including spatial prediction and uncertainty quantification. A more explicit masking approach for handling missing or censored data is also well-established for estimating model parameters in SBI. Wang et al. (2024); Richards et al. (2024); Sainsbury-Dale et al. (2025) include a mask labeling missing or censored data as input to the neural network for parameter inference tasks.
@@
Here, we combine the masking approach of Gloeckler et al. (2024) with the more explicit masking approach of Richards et al. (2024) for the spatial process model setting. The result is a conditional score-based diffusion model within the SDE framework which can quickly and accurately conditionally simulate from high-dimensional predictive distributions. Specifically, we train a neural network that approximates the time-evolving conditional score function, taking as inputs the observed locations, the observed values, the spatial process model parameters, and the diffusion timestep. Importantly, training the neural network only requires unconditional simulation from the spatial process model, which is feasible for a broad class of models. Once trained, the neural network can be used to evaluate the conditional score function in the conditional reverse process, enabling conditional sampling for any observations—regardless of value, number, and location of these observations—and any model parameters without retraining the neural network. Our approach to conditional simulation, which we refer to as neural conditional simulation (NCS), is said to be amortized with respect to these quantities.
@@
We carry out experiments comparing NCS to classical methods for conditional simulation with two spatial processes: (i) a Gaussian process and (ii) a Brown–Resnick process (Brown and Resnick, 1977; Kabluchko et al., 2009), which is a non-Gaussian max-stable process commonly used to model spatial extremes (Huser and Davison, 2014; Engelke et al., 2015; Davison and Huser, 2015). In the former case, exact conditional simulation is both tractable and computationally efficient; this case study serves to validate NCS and verify that it correctly samples from the true predictive distribution. For the latter case, exact conditional simulation is computationally and analytically intractable in moderate to high dimensions (Dombry et al., 2013); the standard approximation involving Markov Chain Monte Carlo (MCMC) has been implemented in the SpatialExtremes package (Ribatet, 2020). We show that NCS is both more accurate and computationally efficient than this standard approximation across a range of observed patterns and spatial process model parameters. From these case studies, we conclude that NCS is a viable methodology for sampling from intractable predictive distributions of spatial processes with both high accuracy and computational efficiency.
@@
The paper is organized as follows. Section 2 provides an overview of score-based diffusion models for generative modeling, and describes how they can be adapted for conditional simulation of spatial processes. In Section 3, we evaluate NCS on simulations from a Brown–Resnick spatial process and show that the NCS-generated simulations are representative of the true predictive distributions via a variety of evaluation metrics. Finally, in Section 4, we discuss the limitations of NCS as well as possible extensions.
@@ 2.1 General Setting
Consider a finite-variance spatial process $h(\cdot)$ dependent on parameter $\bm{\theta}$ and defined on a spatial domain $\mathcal{D}\subset\mathbb{R}^{d}$ (typically $d=2,3)$. Denote by $\mathcal{S}\mathrel{\mathop{:}}=\{\bm{s}_{i}\in\mathcal{D}:i\in\{1,\dots,n\}$ a collection of $n$ locations. Further, let $\mathring{\mathcal{S}}\subset\mathcal{S}$ be the set of spatial locations at which we observe $h(\cdot)$, and let $\mathcal{\tilde{S}}\mathrel{\mathop{:}}=\mathcal{S}\setminus\mathring{\mathcal{S}}$ be the set of spatial locations at which we wish to predict $h(\cdot)$. The task of conditional simulation can be summarized as follows:
Process: $\leavevmode\nobreak\ h(\cdot)\sim\textrm{SpatialProcess}(\bm{\theta}),$ (1a)
Observe: $\leavevmode\nobreak\ \bm{\mathring{x}}_{0}:=h(\mathring{\mathcal{S}}),$ (1b)
Conditionally simulate: $\leavevmode\nobreak\ \tilde{\bm{x}}_{0}\sim p(\tilde{\bm{x}}_{0}\mid\bm{\mathring{x}}_{0};\bm{\theta}),$ (1c)
@@ 2.2 Unconditional Score-Based Diffusion
Here, we first consider the case in which we want to sample $\bm{x}_{0}:=h(\mathcal{S})\sim p(\bm{x}_{0})$, $\bm{x}_{0}\in\mathbb{R}^{\lvert\mathcal{S}\rvert}$, where $\lvert\mathcal{S}\rvert=n$ is the number of locations, via an unconditional score-based diffusion model. In score-based diffusion, one first defines an SDE that diffuses $\bm{x}_{0}$ into a sample $\bm{x}_{T}\sim p(\bm{x}_{T})$, where $p(\bm{x}_{T})$ closely approximates a fixed, predetermined reference distribution $p_{\textrm{ref}}(\bm{x})$ selected for ease of sampling (Pedrotti et al., 2024). Provided the reverse process exists, one can then approximately sample from $p(\bm{x}_{0})$ by first sampling $\bm{x}_{T}\sim p_{\textrm{ref}}(\bm{x})$ and then evolving the sample through the corresponding reverse SDE (Song et al., 2021).
@@
Initial condition: $\leavevmode\nobreak\ \bm{x}_{0}\sim p(\bm{x}_{0}),$ (2a)
Forward evolution: $\leavevmode\nobreak\ \mathrm{d}\bm{x}_{t}=\bm{f}(\bm{x}_{t},t)\mathrm{d}t+g(t)\mathrm{d}\bm{B}_{t},\quad t\in[0,T],$ (2b)
where $\bm{B}_{t}$ is $n$-dimensional Brownian motion, and where the drift and diffusion coefficients $\bm{f}(\cdot,\cdot):\mathbb{R}^{\lvert\mathcal{S}\rvert}\times[0,T]\to\mathbb{R}^{\lvert\mathcal{S}\rvert}$ and $g(\cdot):[0,T]\to\mathbb{R}^{+}$ are user-specified and fixed. Note that time here is artificial and distinct from any temporal dimension in the variables being diffused. The reverse SDE is also a diffusion process given by
Initial condition: $\leavevmode\nobreak\ \bm{x}_{T}\sim p_{\textrm{ref}}(\bm{x}),$ (3a)
Reverse evolution: $\leavevmode\nobreak\ \mathrm{d}\bm{x}_{t}=\big{(}\bm{f}(\bm{x}_{t},t)-g(t)^{2}\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\big{)}\mathrm{d}t+g(t)\mathrm{d}\bm{R}_{t},\quad t\in[0,T],$ (3b)
where $\bm{R}_{t}$ is $n$-dimensional reverse Brownian motion (i.e., Brownian motion that evolves backwards in time) and $p(\bm{x}_{t})$ is the distribution of the samples $\bm{x}_{t}$ from the perturbed target distribution at time $t$ (Anderson, 1982). Note that the drift coefficient in the reverse SDE differs from that in the forward SDE as it includes the score function $\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))$ evaluated at the sample $\bm{x}_{t}$, for time $t\in[0,T]$. If the target distribution $p(\bm{x}_{0})$ is non-Gaussian, $p(\bm{x}_{t})$ cannot be adequately approximated by a Gaussian density for $t$ sufficiently smaller than $T$, and hence the score function is generally analytically or computationally intractable.
Indirect simulation from $p(\bm{x}_{0})$ via the reverse process is possible by replacing the score function in (3b) with an accurate surrogate $\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)$ – a neural network with trainable parameters $\bm{\phi}$ in this case. One way to ensure the surrogate sufficiently approximates the score function for all $\bm{x}_{t}\in\mathbb{R}^{\lvert\mathcal{S}\rvert}$ and $t\in[0,T]$ is by score matching–which involves solving the following minimization problem (Song et al., 2021):
@@
where $\lambda:[0,T]\to\mathbb{R}^{+}$ is a time-varying weighting function, and where the loss function includes an expectation over time to ensure the surrogate $\bm{s}_{\bm{\phi}}(\bm{x}_{t},\cdot)$ learns the score function for any $t\in(0,T)$ in the reverse SDE. For completeness, we provide in Supplementary Materials S1.1 a proof of the second equality in (4) derived from that in Vincent (2011).
@@ 2.3 Conditional Score-Based Diffusion
In this section, we modify the unconditional score-based diffusion approach of Section 2.2 to generate samples from the predictive distribution at $\tilde{\mathcal{S}}$ using observations at $\mathring{\mathcal{S}}$, such that $\mathcal{S}={\tilde{\mathcal{S}}}\cup{\mathring{\mathcal{S}}}$ and $\tilde{\mathcal{S}}\cap\mathring{\mathcal{S}}=\emptyset$. We adapt the approach of Gloeckler et al. (2024) to allow conditional simulation from predictive distributions involving arbitrary partitionings of $\mathcal{S}$ in a spatial setting. We modify the forward and reverse SDEs in (2) and (3), respectively, so that only $\tilde{\bm{x}}_{t}$, the unobserved component of $\bm{x}_{t}=(\tilde{\bm{x}}_{t},\bm{\mathring{x}}_{t})$, is diffused. Specifically, we use the following conditional forward SDE:
Initial condition: $\leavevmode\nobreak\ \tilde{\bm{x}}_{0}\sim p(\tilde{\bm{x}}_{0}\mid\bm{\mathring{x}}_{0}),\quad\bm{\mathring{x}}_{0}\mathrel{\mathop{:}}=h(\mathring{\mathcal{S}}),$ (5a)
Forward evolution: $\leavevmode\nobreak\ \mathrm{d}\tilde{\bm{x}}_{t}=\tilde{\bm{f}}(\tilde{\bm{x}}_{t},t)\mathrm{d}t+g(t)\mathrm{d}\tilde{\bm{B}}_{t},\quad t\in[0,T],$ (5b)
@@
Initial condition: $\leavevmode\nobreak\ \tilde{\bm{x}}_{T}\sim p_{\textrm{ref}}(\tilde{\bm{x}}),\quad\bm{\mathring{x}}_{0}:=h(\mathring{\mathcal{S}}),$ (6a)
Reverse evolution: $\leavevmode\nobreak\ \mathrm{d}\tilde{\bm{x}}_{t}=\big{(}\tilde{\bm{f}}(\tilde{\bm{x}}_{t},t)-g(t)^{2}\nabla_{\tilde{\bm{x}}_{t}}\log(p(\tilde{\bm{x}}_{t}\mid\bm{\mathring{x}}_{0}))\big{)}\mathrm{d}t+g(t)\mathrm{d}\tilde{\bm{R}}_{t},\quad t\in[0,T],$ (6b)
@@
Consider a distribution over the observed locations, $p(\mathring{\mathcal{S}})$, that has positive mass over the set of all possible partitions of $\mathcal{S}$. By the average optimality property for non-negative loss functions (Zammit-Mangion et al., 2025), the modified score matching algorithm is given by:
$\bm{\phi}^{*}$ $=\operatorname*{arg\,min}_{\bm{\phi}}{\mathbb{E}_{t\sim U(0,T)}}\Big{(}\lambda(t){\mathbb{E}_{\mathring{\mathcal{S}}}}{\mathbb{E}_{\bm{\mathring{x}}_{0}}}{\mathbb{E}_{\tilde{\bm{x}}_{t}\mid\bm{\mathring{x}}_{0}}}\Big{[}\left\lVert\tilde{\bm{s}}_{\bm{\phi}}\big{(}{\bm{x}}_{t},\mathring{\mathcal{S}},t\big{)}-\nabla_{\tilde{\bm{x}}_{t}}\log(p(\tilde{\bm{x}}_{t}\mid\bm{\mathring{x}}_{0}))\right\rVert_{2}^{2}\Big{]}\Big{)}$ (7)
@@
$=\operatorname*{arg\,min}_{\bm{\phi}}{\mathbb{E}_{t\sim U(0,T)}}\Big{(}\lambda(t){\mathbb{E}_{\mathring{\mathcal{S}}}}{\mathbb{E}_{\bm{x}_{0}}}{\mathbb{E}_{\tilde{\bm{x}}_{t}\mid\bm{x}_{0}}}\Big{[}\left\lVert\tilde{\bm{s}}_{\bm{\phi}}\big{(}{\bm{x}}_{t},\mathring{\mathcal{S}},t\big{)}-\nabla_{\tilde{\bm{x}}_{t}}\log(p(\tilde{\bm{x}}_{t}\mid\bm{x}_{0}))\right\rVert_{2}^{2}\Big{]}\Big{)},$
where $\tilde{\bm{s}}_{\bm{\phi}}(\cdot,\cdot,\cdot)$ is the output of the surrogate conditional score function corresponding to $\tilde{\mathcal{S}}$, $\bm{x}_{t}=(\bm{\mathring{x}}_{0},\tilde{\bm{x}}_{t})$, and the second line follows from the same argument as that of (4).
Crucially, (7) reveals that one does not need to sample from or have knowledge of the predictive distributions $p(\tilde{\bm{x}}_{0}\mid\bm{\mathring{x}}_{0})$ to train the conditional score-matching surrogate—one only needs to simulate from the user-specified $p(\mathring{\mathcal{S}})$ and the unconditional distribution of the spatial process model before running the forward evaluation (5). Thus, our diffusion model can be used with a wide class of process models that are straightforward to sample from unconditionally, but difficult to sample from conditionally. Note that the output dimension of the surrogate $\tilde{\bm{s}}_{\bm{\phi}}(\cdot,\cdot,\cdot)$ changes with $\lvert\tilde{\mathcal{S}}\rvert$; Section 2.5 describes how to accommodate this in practice. As in Section 2.2, we use a VPSDE, as detailed in the next section.

### v2

# Neural Conditional Simulation for Complex Spatial Processes

@@ Abstract
Keywords: approximate simulation, diffusion model, generative model, likelihood-free inference, simulation-based inference, spatial extremes
@@ 1 Introduction
In recent years, neural networks have been increasingly used to tackle computational challenges in the first step of data analysis: model fitting through parameter inference (Louppe2016; Cramner; Dalmasso2020; Dalmasso2024). These techniques have been applied in various ways to the spatial setting. For instance, Walchessen2024 leverage neural networks to efficiently evaluate the likelihood function of a spatial process model, while Lenzi2021 and SD2024; SD2025 employ neural networks as Bayes estimators for parameter estimation. Both approaches rely on simulation-based inference (SBI) paired with machine learning—the practice of utilizing simulations from a statistical model to train a machine learning model for statistical inference. These methods preclude the need for an explicitly known or computationally tractable likelihood function (Cramner; Dalmasso2020). In contrast, less work has focused on the subsequent—yet equally important—spatial inference task: making predictions and quantifying uncertainty at unobserved locations when the predictive distributions are not available in closed form. Existing methods, which typically do not use SBI and machine learning, include Markov chain Monte Carlo (MCMC) (Gelfand) and variational inference (Jordan1999; Banerjee). However, these approaches often involve a trade-off between computational efficiency and statistical accuracy and usually become impractical when dealing with a large number of observations or prediction locations.
@@
Here, we propose a new method using SBI to train a neural generative model that can simulate from predictive distributions in a spatial setting. Specifically, we employ state-of-the-art diffusion models that are used for image generation (Yang2024), because regularly gridded spatial data can effectively be viewed as image data. These models leverage both forward diffusion and backward evolution to generate samples from a target distribution in a computationally efficient manner. In the forward diffusion process, samples from a complex target distribution are progressively transformed into samples from a simpler reference distribution. This forward process, described by a stochastic differential equation (SDE), is designed to ensure the existence of a corresponding reverse SDE that evolves samples from the reference distribution back into samples from the target distribution (Song2021). This reverse SDE is characterized by a combination of user-defined terms from the forward SDE and an unknown score function, which is approximated using a neural network trained via a specifically designed loss function. Once the neural network is trained, a sample from the complex target distribution can be drawn efficiently by simulating from the reference distribution and applying the learned reversed SDE.
@@
There are several approaches to incorporating conditional information into a diffusion process, many of which originate from the inpainting problem, a topic extensively studied in computer vision (e.g., Lugmayr2022; Quan2024; Simkus2025). Generally, in image inpainting, the objective is to produce realistic, yet diverse images from a partial image. These approaches usually make no assumptions about an underlying statistical model and, consequently, are not designed to approximate any true predictive distributions. In contrast, cardoso2025 adopt an approach that first trains a diffusion model to unconditionally sample from a spatial process and then performs conditional simulation from the predictive distribution via training-free guidance. However, this approach has not been applied to non-Gaussian processes and relies on several approximations during conditional simulation, which can lead to slower and potentially less accurate generation. A few inpainting methods, such as twisted diffusion (Wu2023) and controllable generation (Song2021), have asymptotic guarantees for exact conditional simulation. Yet, these methods have not thus far been validated in a statistically rigorous manner and may not scale easily to high-dimensional data.
@@
A recent contribution by Gloeckler2024 provides a masking approach to incorporate conditional information into the score-based diffusion process that is designed to accurately simulate from conditional distributions of parameters and data including posterior distributions. However, their approach was not designed for high-dimensional conditioning, including spatial prediction and uncertainty quantification. A more explicit masking approach for handling missing or censored data is also well-established for estimating model parameters in SBI. Wang2024, Richards2024, and SD2025 include a mask labeling missing or censored data as input to the neural network for parameter inference tasks.
@@
Here, we combine the masking approach of Gloeckler2024 with the more explicit masking approach of Richards2024 for the spatial process model setting. The result is a conditional score-based diffusion model within the SDE framework which can quickly and accurately conditionally simulate from high-dimensional predictive distributions. Specifically, we train a neural network that approximates the time-evolving conditional score function, taking as inputs the observed locations, the observed values, the spatial process model parameters, and the diffusion timestep. Importantly, training the neural network only requires unconditional simulation from the spatial process model, which is feasible for a broad class of models. Once trained, the neural network can be used to evaluate the conditional score function in the conditional reverse process, enabling conditional sampling for any observations—regardless of value, number, and location of these observations—and any model parameters without retraining the neural network. Our approach to conditional simulation, which we refer to as neural conditional simulation (NCS), is said to be amortized with respect to these quantities.
@@
We carry out experiments comparing NCS to classical methods for conditional simulation with two spatial processes: (i) a Gaussian process and (ii) a Brown–Resnick process (BrownResnick1977; Kabluchko2009), which is a non-Gaussian max-stable process commonly used to model spatial extremes (Huser2014; Engelke2015; Davison2015). In the former case, exact conditional simulation is both tractable and computationally efficient; this case study serves to validate NCS and verify that it correctly samples from the true predictive distribution. For the latter case, exact conditional simulation is computationally and analytically intractable in moderate to high dimensions (Dombry2012); the standard approximation involving Markov Chain Monte Carlo (MCMC) has been implemented in the SpatialExtremes package (SpatialExtremes). We show that NCS is both more accurate and computationally efficient than this standard approximation across a range of observed patterns and spatial process model parameters. From these case studies, we conclude that NCS is a viable methodology for sampling from intractable predictive distributions of spatial processes with both high accuracy and computational efficiency.
@@
The paper is organized as follows. Section 2 provides an overview of score-based diffusion models for generative modeling, and describes how they can be adapted for conditional simulation of spatial processes. In Section 3, we evaluate NCS on simulations from a Brown–Resnick spatial process and show that the NCS-generated simulations are representative of the true predictive distributions via a variety of validation metrics. Finally, in Section 4, we discuss the limitations of NCS as well as possible extensions.
@@ 2.1 General Setting
Consider a finite-variance spatial process $h(\cdot)$ dependent on parameter $\bm{\theta}$ and defined on a spatial domain $\mathcal{D}\subset\mathbb{R}^{d}$ (typically $d=2,3)$. Denote by $\mathcal{S}\mathrel{\mathop{:}}=\{\bm{s}_{i}\in\mathcal{D}:i\in[n]\}$ a collection of $n$ locations, where $[n]$ is shorthand notation for the set $\{1,\dots,n\}$. Further, let $\mathring{\mathcal{S}}\subset\mathcal{S}$ be the set of spatial locations at which we observe $h(\cdot)$, and let $\mathcal{\tilde{S}}\mathrel{\mathop{:}}=\mathcal{S}\setminus\mathring{\mathcal{S}}$ be the set of spatial locations at which we wish to predict $h(\cdot)$. The task of conditional simulation can be summarized as follows:
Process: $\penalty 10000\ h(\cdot)\sim\textrm{SpatialProcess}(\bm{\theta}),$ (1a)
Observe: $\penalty 10000\ \bm{\mathring{x}}_{0}:=h(\mathring{\mathcal{S}}),$ (1b)
Conditionally simulate: $\penalty 10000\ \tilde{\bm{x}}_{0}\sim p(\tilde{\bm{x}}_{0}\mid\bm{\mathring{x}}_{0};\bm{\theta}),$ (1c)
@@ 2.2 Unconditional Score-Based Diffusion
Here, we first consider the case in which we want to sample $\bm{x}_{0}:=h(\mathcal{S})\sim p(\bm{x}_{0})$, $\bm{x}_{0}\in\mathbb{R}^{\lvert\mathcal{S}\rvert}$}, where $\lvert\mathcal{S}\rvert=n$ is the number of locations, via an unconditional score-based diffusion model. In score-based diffusion, one first defines an SDE that diffuses $\bm{x}_{0}$ into a sample $\bm{x}_{T}\sim p(\bm{x}_{T})$, where $p(\bm{x}_{T})$ closely approximates a fixed, predetermined reference distribution $p_{\textrm{ref}}(\bm{x})$ selected for ease of sampling (Pedrotti2024). Provided the reverse process exists, one can then approximately sample from $p(\bm{x}_{0})$ by first sampling $\bm{x}_{T}\sim p_{\textrm{ref}}(\bm{x})$ and then evolving the sample through the corresponding reverse SDE (Song2021).
@@
Initial condition: $\penalty 10000\ \bm{x}_{0}\sim p(\bm{x}_{0}),$ (2a)
Forward evolution: $\penalty 10000\ \mathrm{d}\bm{x}_{t}=\bm{f}(\bm{x}_{t},t)\mathrm{d}t+g(t)\mathrm{d}\bm{B}_{t},\quad t\in[0,T],$ (2b)
where $\bm{B}_{t}$ is an $n$-dimensional Brownian motion, and where the drift and diffusion coefficients $\bm{f}(\cdot,\cdot):\mathbb{R}^{\lvert\mathcal{S}\rvert}\times[0,T]\to\mathbb{R}^{\lvert\mathcal{S}\rvert}$ and $g(\cdot):[0,T]\to\mathbb{R}^{+}$ are user-specified and fixed. Note that time here is artificial and distinct from any temporal dimension in the variables being diffused. The reverse SDE is also a diffusion process given by
Initial condition: $\penalty 10000\ \bm{x}_{T}\sim p_{\textrm{ref}}(\bm{x}),$ (3a)
Reverse evolution: $\penalty 10000\ \mathrm{d}\bm{x}_{t}=\big(\bm{f}(\bm{x}_{t},t)-g(t)^{2}\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))\big)\mathrm{d}t+g(t)\mathrm{d}\bm{R}_{t},\quad t\in[0,T],$ (3b)
where $\bm{R}_{t}$ is an $n$-dimensional reverse Brownian motion (i.e., a Brownian motion that evolves backwards in time) and $p(\bm{x}_{t})$ is the distribution of the samples $\bm{x}_{t}$ from the perturbed target distribution at time $t$ (Anderson1982). Note that the drift coefficient in the reverse SDE differs from that in the forward SDE as it includes the score function $\nabla_{\bm{x}_{t}}\log(p(\bm{x}_{t}))$ evaluated at the sample $\bm{x}_{t}$, for time $t\in[0,T]$. If the target distribution $p(\bm{x}_{0})$ is non-Gaussian, $p(\bm{x}_{t})$ cannot be adequately approximated by a Gaussian density for $t$ sufficiently smaller than $T$, and hence the score function is generally analytically or computationally intractable.
Indirect simulation from $p(\bm{x}_{0})$ via the reverse process is possible by replacing the score function in (3b) with an accurate surrogate $\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)$—a neural network with trainable parameters $\bm{\phi}$ in this case. One way to ensure the surrogate sufficiently approximates the score function for all $\bm{x}_{t}\in\mathbb{R}^{\lvert\mathcal{S}\rvert}$ and $t\in[0,T]$ is by score matching, which involves solving the following minimization problem (Song2021):
@@
where $\lambda:[0,T]\to\mathbb{R}^{+}$ is a time-varying weighting function, and where the loss function includes an expectation over time to ensure the surrogate $\bm{s}_{\bm{\phi}}(\bm{x}_{t},\cdot)$ learns the score function for any $t\in(0,T)$ in the reverse SDE. For completeness, we provide in Supplementary Materials S1.1 a proof of the second equality in (4) derived from that in Vincent2011.
@@ 2.3 Conditional Score-Based Diffusion
In this section, we modify the unconditional score-based diffusion approach of Section 2.2 to generate samples from the predictive distribution at $\tilde{\mathcal{S}}$ using observations at $\mathring{\mathcal{S}}$, such that $\mathcal{S}={\tilde{\mathcal{S}}}\cup{\mathring{\mathcal{S}}}$ and $\tilde{\mathcal{S}}\cap\mathring{\mathcal{S}}=\emptyset$. We adapt the approach of Gloeckler2024 to allow conditional simulation from predictive distributions involving arbitrary partitionings of $\mathcal{S}$ in a spatial setting. We modify the forward and reverse SDEs in (2) and (3), respectively, so that only $\tilde{\bm{x}}_{t}$, the unobserved component of $\bm{x}_{t}=(\tilde{\bm{x}}^{\prime}_{t},\bm{\mathring{x}}^{\prime}_{t})^{\prime}$, is diffused. Specifically, we use the following conditional forward SDE:
Initial condition: $\penalty 10000\ \tilde{\bm{x}}_{0}\sim p(\tilde{\bm{x}}_{0}\mid\bm{\mathring{x}}_{0}),\quad\bm{\mathring{x}}_{0}\mathrel{\mathop{:}}=h(\mathring{\mathcal{S}}),$ (5a)
Forward evolution: $\penalty 10000\ \mathrm{d}\tilde{\bm{x}}_{t}=\tilde{\bm{f}}(\tilde{\bm{x}}_{t},t)\mathrm{d}t+g(t)\mathrm{d}\tilde{\bm{B}}_{t},\quad t\in[0,T],$ (5b)
@@
Initial condition: $\penalty 10000\ \tilde{\bm{x}}_{T}\sim p_{\textrm{ref}}(\tilde{\bm{x}}),\quad\bm{\mathring{x}}_{0}:=h(\mathring{\mathcal{S}}),$ (6a)
Reverse evolution: $\penalty 10000\ \mathrm{d}\tilde{\bm{x}}_{t}=\big(\tilde{\bm{f}}(\tilde{\bm{x}}_{t},t)-g(t)^{2}\nabla_{\tilde{\bm{x}}_{t}}\log(p(\tilde{\bm{x}}_{t}\mid\bm{\mathring{x}}_{0}))\big)\mathrm{d}t+g(t)\mathrm{d}\tilde{\bm{R}}_{t},\quad t\in[0,T],$ (6b)
@@
Consider a distribution over the observed locations, $p(\mathring{\mathcal{S}})$, that has positive mass over the set of all possible partitions of $\mathcal{S}$. By the average optimality property for non-negative loss functions (Zammit2024), the modified score matching algorithm is given by:
$\bm{\phi}^{*}$ $=\operatorname*{arg\,min}_{\bm{\phi}}{\mathbb{E}_{t\sim U(0,T)}}\Big(\lambda(t){\mathbb{E}_{\mathring{\mathcal{S}}}}{\mathbb{E}_{\bm{\mathring{x}}_{0}}}{\mathbb{E}_{\tilde{\bm{x}}_{t}\mid\bm{\mathring{x}}_{0}}}\Big[\left\lVert\tilde{\bm{s}}_{\bm{\phi}}\big({\bm{x}}_{t},\mathring{\mathcal{S}},t\big)-\nabla_{\tilde{\bm{x}}_{t}}\log(p(\tilde{\bm{x}}_{t}\mid\bm{\mathring{x}}_{0}))\right\rVert_{2}^{2}\Big]\Big)$
@@
$=\operatorname*{arg\,min}_{\bm{\phi}}{\mathbb{E}_{t\sim U(0,T)}}\Big(\lambda(t){\mathbb{E}_{\mathring{\mathcal{S}}}}{\mathbb{E}_{\bm{x}_{0}}}{\mathbb{E}_{\tilde{\bm{x}}_{t}\mid\bm{x}_{0}}}\Big[\left\lVert\tilde{\bm{s}}_{\bm{\phi}}\big({\bm{x}}_{t},\mathring{\mathcal{S}},t\big)-\nabla_{\tilde{\bm{x}}_{t}}\log(p(\tilde{\bm{x}}_{t}\mid\bm{x}_{0}))\right\rVert_{2}^{2}\Big]\Big),$ (7a)
where $\tilde{\bm{s}}_{\bm{\phi}}(\cdot,\cdot,\cdot)$ is the output of the surrogate conditional score function corresponding to $\tilde{\mathcal{S}}$, $\bm{x}_{t}=(\bm{\mathring{x}}_{0}^{\prime},\tilde{\bm{x}}_{t}^{\prime})^{\prime}$, and the second line follows from the same argument as that of (4).
Crucially, (7a) reveals that one does not need to sample from or have knowledge of the predictive distributions $p(\tilde{\bm{x}}_{0}\mid\bm{\mathring{x}}_{0})$ to train the conditional score-matching surrogate—one only needs to simulate from the user-specified $p(\mathring{\mathcal{S}})$ and the unconditional distribution of the spatial process model before running the forward evolution (5). Thus, our diffusion model can be used with a wide class of process models that are straightforward to sample from unconditionally, but difficult to sample from conditionally. Note that the output dimension of the surrogate $\tilde{\bm{s}}_{\bm{\phi}}(\cdot,\cdot,\cdot)$ changes with $\lvert\tilde{\mathcal{S}}\rvert$; Section 2.5 describes how to accommodate this in practice. As in Section 2.2, we use a VPSDE, as detailed in the next section.